write the equation of the perpendicular bisector of the segment joining A(-5,4) and B(3,-6)
Answer by benazir.sj@gmail.com(109)


write the equation of the perpendicular bisector of the segment joining A(-5,4) and B(3,-6)
solution:
Bisector passes through midpoint:
midpoint of (-5,4) and (3,-6) = ((-5+4)/2, (3+(-6))/2)
=(-1/2,-3/2)
Perpendicular bisector has slope that is negative reciprocal of line segment joining points (-5,4) and (3,-6)
Slope of line segment = y2-y1/x2-x1
=-6-4/3+5
=-10/8 = -5/4
negative reciprocal =4/5
Perpendicular bisector has slope = 4/5 and passes through point (-1/2,-3/2)
Use point-slope form:
y + 3/2 = 4/5 (x +1/2)
y+3/2 = 4/5x+ 2/5
y= 4/5x +2/5 -3/2
y = 4/5x-11/10